# Word-length curve counting on the once-punctured torus

### v2

@@ Abstract.
We classify closed curves on a once-punctured torus with a single self-intersection from a combinatorial perspective. We determine the number of closed curves with given word-length and with zero, one, and arbitrary self-intersections.
@@
DF is supported by the Luxembourg National Research Fund PRIDE17/1224660/GPS. ML is supported by the Luxembourg National Research Fund OPEN grant O19/13865598.
@@ I.   Introduction
Figure 1. A once-punctured torus.
@@
where $c$ runs over the set of closed curves on $\Sigma_{1,1}$ in the free homotopy class $\gamma$. Curves with 00 self-intersection are commonly referred to as simple curves.
The fundamental group $\pi_{1}(\Sigma_{1,1})$ is a free group of rank $2$. By choosing a pair $(a,b)$ of generators for the fundamental group (see Figure 1 for an example), we can identify $\pi_{1}(\Sigma_{1,1})$ with the free group $\mathbf{F}_{2}$ generated by $a$ and $b$. A word is a finite sequence $(x_{i})_{i=1}^{n}$ such that $x_{i}\in\{a,b,a^{-1},b^{-1}\}$ for all $1\leq i\leq n$. A circular word can be seen as a word written on a circle, or more formally, a circular word is an equivalence class of words where two words are equivalent if they differ by a circular shift. A word $(x_{i})_{i}^{n}$ is called reduced if $x_{i+1}\neq x_{i}^{-1}$ for all $i=1,\dots,n-1$, and $x_{1}\neq x_{n}^{-1}$. A circular word is reduced if one (or all) of its representatives are reduced. In this note, we consider only reduced (circular) words. Words can be seen as elements in $\mathbf{F}_{2}$, and circular words can be seen as conjugacy classes in $\mathbf{F}_{2}$. Curves in $\mathscr{C}(\Sigma_{1,1})$ are in one-to-one correspondence with reduced circular words. We say that a circular word is primitive if none of its representatives is a proper power of another word. Primitive curves correspond to primitive reduced circular words.
@@ I.1   Results
The following result is classic. However, we shall give a new proof of combinatorial nature, whose ingredients are useful later.
@@
A simple multicurve is a formal sum of pairwise non-homotopic primitive simple curves with positive integer coefficients. On $\Sigma_{1,1}$, the set of multicurves is nothing but $\mathscr{C}(\Sigma_{1,1})$. The last result allows us to count multicurves on $\Sigma_{1,1}$, which the authors found missing in the literature.
@@
    $a^{n_{1}}b\cdots a^{n_{k}}b$, where $[n_{1},\ldots,n_{k}]$ satisfies that exists an $m\in\mathbb{Z}_{\geq 1}$ such that for all $i\in\{1,\ldots,k\}$, $n_{i}\in\{m,m+1\}$ and it is a necklace with $2$-variation (see Definition III.4), or
@@
The non-necessarily primitive case is just the sum of the number of primitive ones plus the number of primitive simple ones with half the length (as these will square to a single self-intersection). The classification allows us to again do the exact counting for any $L$, via using the techniques from the proofs of the simple case.
@@
Remark I.6.
It was brought to the attention of the authors, after finding this counting as a consequence of the classification, that some strategy of the counting is shared at Moira Chas’ Preprint [Cha, Proposition 3.2].
@@
There are $4$ primtive curves of length $1$, $8$ of length $2$, and for any $L\in\mathbb{Z}_{\geq 3}$, we have formula
[…]
@@
[…]
@@
=|$=|\{\text{aperiodic necklaces with $L$ beads and }3\text{ colors}\}|+\delta_{\{1,2\}}(L),$
@@ I.2   Related work
One can also count curves under more topological constraints. For example, one may restrict consideration to curves with no self-intersection, or a given number of self-intersections. Efforts in this direction (see for example [Ree81, BS85, McSR95, Riv01]) reached their greatest heights with Mirzakhani’s landbreaking work [Mir08]. As a consequence of her findings, there exist explicit constants $C_{g,n}>0$ depending only on $g$ and $n$ and $B_{X}>0$ depending only on the hyperbolic metric $X$ such that
@@
This result was extended to the non-simple case by Mirzakhani twelve years later [Mir]: for any $k\in\mathbb{Z}_{\geq 0}$ there exists a constant $C_{g,n;k}$, such that (I.2) remains valid when replacing “$i(\gamma)=0$” by “$i(\gamma)=k$” and $C_{g,n}$ by $C_{g,n;k}$.
@@ II.1   Curves as necklaces
We use “()($\dots$)( … )” to denote sequences, and use “[]delimited-[][$\dots$][ … ]” to denote necklaces. For instance, $(1,2,3)$, $(2,3,1)$, and $(3,1,2)$ are different sequences but represent the same necklace $[1,2,3]$.
@@
A necklace $[n_{i}]_{i}$ has small variation if, for any $s\in\mathbb{Z}_{\geq 1}$, sums of $s$ consecutive elements never differ by more than ±$\pm 1$. In symbols, this means
@@
Let $w=[n_{i}]_{i}$ be a non-constant necklace. If (II.1) holds for $s=1$, then there exists $m\in\mathbb{Z}$ such that $n_{i}\in\{m,m+1\}$ for all index $i$. If (II.1) holds for $s=2$ too, then $m$ (resp. $m+1$) cannot appear consecutively if $|w|_{m}\leq|w|_{m+1}$ (resp. $|w|_{m+1}\leq|w|_{m}$).
@@
    $ab^{n_{1}}ab^{n_{2}}\cdots ab^{n_{r}}$, where $[n_{1},\dots,n_{r}]$ has small variation.
@@
We say that a simple closed curve on $\Sigma_{1,1}$ has general type if it falls into the third case described in Theorem II.5. For such a curve, we associate it with the necklace $[n_{1},\dots,n_{r}]$, which we will refer to as its exponent necklace. See Figure 2(a) and 2(b) for an example.
Write $M(L)$ for the set of closed multicurves of general type on $\Sigma_{1,1}$ of length $L$, and $N(L)$ for the set of necklaces $[n_{1},\dots,n_{r}]$ such that $r+\sum_{i}n_{i}=L$.
@@
For $L$ odd, $M(L)$ and $N(L)$ are in eight-to-one correspondence. For $L$ even, we have $|M(L)|=8|N(L)|-4$.
@@
First, it is not hard to check that a power of a necklace (the square of $[1,2,3]$ is $[1,2,3,1,2,3]$) with small variation has small variation. Thus every power of a (primitive) simple curve of general type is of general type. The eight-to-one correspondence in the case where $L$ is odd arises from the $8$ possible renamings of the generators. This correspondence breaks down when $L$ is even due to the fact that the necklace $[1,\dots,1]$ corresponds to only $4$ curves (for example, $ab\cdots ab$ and $ba\cdots ba$ give the same curve). ∎
@@ II.2   Necklaces with small variation
Let $w$ be a necklace with small variation. Recall that we write $|w|_{m}$ for the number of occurrences of $m$ in $w$. We define a new necklace $A(w)$ as follows. By Remark II.3, there exists $m\in\mathbb{Z}_{\geq 1}$ such that $n_{i}\in\{m,m+1\}$ for all $i$. If $|w|_{m}\leq|w|_{m+1}$ (resp. $|w|_{m+1}\leq|w|_{m}$), we define $A(w)$ to be the necklace obtained by removing all the $m$’s (resp. $(m+1)$’s) from $w$ and replacing each run of $m+1$ (resp. $m$) by the length of the run. For example, $A[4,5,5,4,5,5,5,4,5,5]=[2,3,2]$ and $A[4,5,5,5,4,5,4,5,5]=[3,1,2]$. Note that if $|w|_{m}=|w|_{m+1}$, then $A(w)$ is a constant necklace $[1,\dots,1]$ of size $|w|_{m}=|w|_{m+1}$.
@@ II.3   Counting
For any positive integer $L$, there are exactly $4(L-1)$ simple multicurves of general type and length $L$.
@@
where $L\in\mathbb{Z}_{\geq 1}$ is given, and $x,m\in\mathbb{Z}_{\geq 1}$, $y\in\mathbb{Z}_{\geq 0}$ are unknown. Then, there are exactly $L/2$ solutions $(x,y,m)$ if $L$ is even, and $(L-1)/2$ solutions if $L$ is odd.
@@
By Proposition II.11, $m(n)=4(n-1)$. Define also
@@
For the second part, by Proposition II.11, we know that the number of powers of simple closed curves of general type with word length exactly $L$ is $M(L)=4(L-1)$. Hence,
@@ III.1   Characterization
Let $\omega=x_{1}\cdots x_{n}$, where $x_{i}\in\{a,b,a^{-1},b^{-1}\}$ be a reduced word. Consider all of its circular shifts $\{\omega_{i}\}_{i=1}^{n}$ where $\omega_{i}=x_{i}x_{i+1}\cdots x_{n}x_{1}x_{2}\cdots x_{i-1}$ and the cyclic lexicographic ordering induced by $a<b<a^{-1}<b^{-1}$: two words $\alpha=x_{1}\cdots x_{n}$ and $\beta=y_{1}\cdots y_{n}$ satisfy $\alpha<\beta$ if either $x_{1}<y_{1}$, or for some $1<t\leq n$, we have $x_{i}=y_{i}$ for $11\leq i\leq t-$ 1, and $x_{t}<y_{t}$ under the new ordering obtained by cyclically shifting the original one until it starts with $x_{t-1}^{-1}$.
@@
Figure 4. Surgery on a word of the form $ab^{-1}a^{-i}b\omega$ with $i\geq 2$.
@@
Hence, we can restrict to the case where the changes of sign come from subwords of the kind $ab^{-1}a^{-1}b$ (up to renaming $a$ for $a^{-1}$ or $b$ for $b^{-1}$). Hence, take a reduced word $ab^{-1}a^{-1}b\omega$ with $\omega$ a subword such that the exponents on $b$ are ±$\pm 1$. Note that if $\omega$ starts with an $a$, up to homotopy, one finds the first situation in Figure 5, meaning that the surface is divided into two regions with the startpoint of the curve in one of them and the continuation in the other one. Thus, the curve will intersect at some point the arcs given by the commutator $ab^{-1}a^{-1}b$, and so $i(\omega)\leq i(ab^{-1}a^{-1}b\omega)-1$. Therefore, $\omega$ has to represent a simple word starting with $a$. Moreover, $\omega$ has to be of the form $a^{n_{1}}b\cdots a^{n_{k}}b$, as if it was of the form $a^{n_{1}}b^{-1}\cdots a^{n_{k}}b^{-1}$ the curve $ab^{-1}a^{-1}b\omega$ would contain the subword $ba^{n_{1}}b^{-1}a^{n_{2}}$ and the only such a word with self-intersection one is $bab^{-1}a$ as proved above, or if $\omega=ab^{-1}a^{-1}b$, the word would not be primitive.
@@
Note now that the maps $\alpha_{m},\tilde{\alpha}_{m}$ defined in (II.2) fix the set of conjugacy classes $\{[ab^{-1}a^{-1}b],[(ab^{-1}a^{-1}b)^{-1}]\}$. Hence, considering a reduced circular word of the kind $ab^{-1}a^{-1}b\omega$, such that it has a single self-intersection and such that $\omega=a^{n_{1}}b\cdots a^{n_{k}}b$ with $n_{i}\in\{m,m+1\}$ for some $m\neq 0,1$ represents a primitive simple word, we can apply the reduction $\alpha_{m}(ab^{-1}a^{-1}b\omega)=\omega^{\prime}\alpha_{m}(\omega)$ if |$|\{i\mid n_{i}=m\}|>|\{i\mid n_{i}=m+1\}|$ (or with $\tilde{\alpha}_{m}$ otherwise), where $\omega^{\prime}$ is a cyclically reduced representative of the conjugacy classes of the commutators of $a,b$. As we are assuming that $\omega$ represents a simple word, applying finitely many times maps of the family $\alpha_{m}$, $\tilde{\alpha}_{m}$, this reduction acts as $A$ on the necklace $[n_{1},\ldots,n_{k}]$ defined in Figure 2(c), and it will reduce to a word of the kind $A^{n}(\omega)=a^{n^{\prime}}b$ up to renaming the generators. Therefore, by inversing $A$ with the appropriate $B_{m}$’s, and given Lemma III.3, after renaming properly the generators such that the representative of the commutator is of the form $ab^{-1}a^{-1}b$, we find that there is a unique permutation of every simple word $\omega$ of the form $a^{n_{1}}b\cdots a^{n_{k}}b$ or $a^{-n_{1}}b\cdots a^{-n_{k}}b$ making $ab^{-1}a^{-1}b\omega$ a self-intersection one curve.
Moreover, applying this reduction to Lemma III.3, one sees that, in order to obtain a self-intersection one curve from a simple one of the form $a^{n_{1}}b\cdots a^{n_{k}}b$ with $n_{i}\in\{m,m+1\}$ for some $m\neq 0,-1$, the commutator $ab^{-1}a^{-1}b$, has to be inserted before a block $a^{n_{i}}b$ such that $n_{i}=m+1$, $n_{i-1}=m$ and maximizing as much as possible the beginning of the chain $\{\sum_{j=1}^{s}|n_{i+\epsilon j}|\}_{s\geq 1}$ with $\epsilon=1$ for positive $m$ and $-1$ for negative (e.g. see Figure 6).
@@
  • $a^{n_{1}}b\cdots a^{n_{k}}b$ where the exponent necklace $[n_{1},\dots,n_{k}]$ has $2$-variation.
@@
Let us now move to the case where we have a word of the form $a^{n_{1}}b\cdots a^{n_{k}}b$ with all $n_{i}\in\{m,m+1\}$ for some $m\in\mathbb{Z}_{\geq 1}$. We want to prove that it has self-intersection one if and only if the necklace of positive integers $[n_{1},\ldots,n_{k}]$ has $2$-variation.
It is enough to see that in this case there is a $1:1$ correspondence between classes of linking pairs in [CL87] algorithm and essential pairs of blocks in the necklace $[n_{1},\ldots,n_{k}]$. We will start by constructing a class of linking pairs in the word given an essential pair of blocks in the associated necklace.

### v3

# Single self-insterction words on the once-punctured torus and their counting

@@ Abstract.
We classify closed curves on a once-punctured torus with a single self-intersection from a combinatorial perspective. We then use the classification to determine the number of closed curves with given word-length and with zero, one, and arbitrary self-intersections.
@@
DF is supported by the Luxembourg National Research Fund PRIDE17/1224660/GPS and acknowledges support by the FEDER/AEI/MICINN grant PID2021-125625NB-I00 “Estructuras y Desigualdades Geométricas Universales” and the AGAUR grant 2021-SGR-01015.
ML is supported by the Luxembourg National Research Fund OPEN grant O19/13865598.
@@ I.  Introduction
Figure 1. A once-punctured torus
@@
where $c$ runs over the set of closed curves on $\Sigma_{1,1}$ in the free homotopy class $\gamma$. Curves with 0 self-intersection are commonly referred to as simple curves.
The fundamental group $\pi_{1}(\Sigma_{1,1})$ is a free group of rank $2$. By choosing a pair $(a,b)$ of generators for the fundamental group (see Figure 1 for an example), we can identify $\pi_{1}(\Sigma_{1,1})$ with the free group $\mathbf{F}_{2}$ generated by $a$ and $b$. A word is a finite sequence $(x_{i})_{i=1}^{n}$ such that $x_{i}\in\{a,b,a^{-1},b^{-1}\}$ for all $1\leq i\leq n$. A circular word can be seen as a word written on a circle, or more formally, an equivalence class of words where two words are equivalent if they differ by a circular shift. A word $(x_{i})_{i}^{n}$ is called reduced if $x_{i+1}\neq x_{i}^{-1}$ for all $i=1,\dots,n-1$, and $x_{1}\neq x_{n}^{-1}$. A circular word is reduced if one (or all) of its representatives are reduced. In this note, we consider only reduced (circular) words. Words can be seen as elements in $\mathbf{F}_{2}$, and circular words can be seen as conjugacy classes in $\mathbf{F}_{2}$. Curves in $\mathscr{C}(\Sigma_{1,1})$ are in one-to-one correspondence with reduced circular words. We say that a circular word is primitive if none of its representatives is a proper power of another word. Primitive curves correspond to primitive reduced circular words.
@@ I.1  Results
The results will appear in the order they will be proved during the article. The first one is a well-known result as a consequence of the homotopy-homology correspondence on the once-punctured torus. We will give a new proof of combinatorial nature, whose ingredients are useful later to prove a higher intersection case.
@@
A simple multicurve is a formal sum of pairwise non-homotopic disjoint simple curves with positive integer coefficients. On $\Sigma_{1,1}$, the set of multicurves is nothing but $\mathscr{C}(\Sigma_{1,1})$. The last result allows us to count multicurves on $\Sigma_{1,1}$, which the authors found missing in the literature.
@@
    $a^{n_{1}}b\cdots a^{n_{k}}b$, where $n_{1},\ldots,n_{k}]$ satisfies that exists an $m\in\mathbb{Z}_{\geq 1}$ such that for all $i\in\{1,\ldots,k\}$, $n_{i}\in\{m,m+1\}$ and it is a necklace with $2$-variation (see Definition III.4), or
@@
The non-necessarily primitive case is just the sum of the number of primitive ones plus the number of primitive simple ones with half the length - as these will square to a single self-intersection. The classification allows us again to do the exact count for any $L$, using techniques from the proofs of the simple case. As in the case of simple closed curves, this counting result can already be found in the literature, more specifically in Moira Chas’ Preprint [Cha, Proposition 3.2]. This new combinatorial proof is longer than the existing one due to the use of new combinatorial objects and the need to check further details.
@@
There are $4$ primtive curves of length $1$, 88 of length $2$, and for any $L\in\mathbb{Z}_{\geq 3}$, we have formula
$|\{\gamma\in\mathscr{P}\mathscr{C}^{*}(\Sigma_{1,1})\mid\ell_{\omega}(\gamma)=L\}|=\frac{1}{L}\sum_{d|L}\mu(d)\,3^{L/d},$
@@
$|\{\gamma\in\mathscr{C}^{*}(\Sigma_{1,1})\mid\ell_{\omega}(\gamma)=L\}|=\frac{1}{L}\sum_{d\mid L}\varphi(d)\,3^{L/d}+\frac{3+(-1)^{L}}{2}.$
@@
$=|\{\text{aperiodic necklaces with $L$ beads and }3\text{ colors}\}|+\delta_{\{1,2\}}(L),$
@@ I.2  Related work
One can also count curves under more topological constraints. For example, one may restrict consideration to curves with no self-intersection, or a given number of self-intersections. Efforts in this direction (see for example [Ree81, BS85, McSR95, Riv01]) reached their greatest heights with Mirzakhani’s groundbreaking work [Mir08]. As a consequence of her findings, there exist explicit constants $C_{g,n}>0$ depending only on $g$ and $n$ and $B_{X}>0$ depending only on the hyperbolic metric $X$ such that
@@
This result was extended to the non-simple case by Mirzakhani twelve years later [Mir]: for any $k\in\mathbb{Z}_{\geq 0}$ there exists a constant $C_{g,n;k}$, such that (I.2) remains valid when replacing $i(\gamma)=0$” by $i(\gamma)=k$” and $C_{g,n}$ by $C_{g,n;k}$.
@@ II.1  Curves as necklaces
We use $(\dots)$” to denote sequences, and use $\dots$]” to denote necklaces. For instance, $(1,2,3)$, $(2,3,1)$, and $(3,1,2)$ are different sequences but represent the same necklace $[1,2,3]$.
@@
A necklace $n_{i}]_{i}$ has small variation if, for any $s\in\mathbb{Z}_{\geq 1}$, sums of $s$ consecutive elements never differ by more than $\pm 1$. In symbols, this means
@@
Let $w=[n_{i}]_{i}$ be a non-constant necklace. If (II.1) holds for $s=1$, then there exists $m\in\mathbb{Z}$ such that $n_{i}\in\{m,m+1\}$ for all index ii. If (II.1) holds for $s=2$ too, then $m$ (resp. $m+1$) cannot appear consecutively if $|w|_{m}\leq|w|_{m+1}$ (resp. $|w|_{m+1}\leq|w|_{m}$).
@@
    $ab^{n_{1}}ab^{n_{2}}\cdots ab^{n_{r}}$, where $n_{1},\dots,n_{r}]$ has small variation.
@@
We say that a simple closed curve on $\Sigma_{1,1}$ has general type if it falls into the third case described in Theorem II.5. For such a curve, we associate it with the necklace $n_{1},\dots,n_{r}]$, which we will refer to as its exponent necklace. See Figure 2(a) and 2(b) for an example.
Write $M(L)$ for the set of closed multicurves of general type on $\Sigma_{1,1}$ of length $L$, and $N(L)$ for the set of necklaces $n_{1},\dots,n_{r}]$ such that $r+\sum_{i}n_{i}=L$.
@@
For $L$ odd, $M(L)$ and $N(L)$ are in eight-to-one correspondence. For $L$ even, we have |M(L)|=8|N(L)|4|M(L)|=8|N(L)|-4.
@@
First, it is not hard to check that a power of a necklace (the square of $[1,2,3]$ is $[1,2,3,1,2,3]$) with small variation has small variation. Thus every power of a (primitive) simple curve of general type is of general type. The eight-to-one correspondence in the case where $L$ is odd arises from the $8$ possible renamings of the generators. This correspondence breaks down when $L$ is even due to the fact that the necklace $1,\dots,1]$ corresponds to only $4$ curves (for example, $ab\cdots ab$ and $ba\cdots ba$ give the same curve). ∎
@@ II.2  Necklaces with small variation
Let $w$ be a necklace with small variation. Recall that we write $|w|_{m}$ for the number of occurrences of $m$ in $w$. We define a new necklace $A(w)$ as follows. By Remark II.3, there exists $m\in\mathbb{Z}_{\geq 1}$ such that $n_{i}\in\{m,m+1\}$ for all $i$. If $|w|_{m}\leq|w|_{m+1}$ (resp. $|w|_{m+1}\leq|w|_{m}$), we define $A(w)$ to be the necklace obtained by removing all the $m$’s (resp. $(m+1)$’s) from $w$ and replacing each run of $m+1$ (resp. $m$) by the length of the run. For example, $A[4,5,5,4,5,5,5,4,5,5]=[2,3,2]$ and $A[4,5,5,5,4,5,4,5,5]=[3,1,2]$. Note that if $|w|_{m}=|w|_{m+1}$, then $A(w)$ is a constant necklace $1,\dots,1]$ of size $|w|_{m}=|w|_{m+1}$.
@@ II.3  Counting
For any positive integer $L$, there are exactly 4(L1)4(L-1) simple multicurves of general type and length $L$.
@@
where $L\in\mathbb{Z}_{\geq 1}$ is given, and $x,m\in\mathbb{Z}_{\geq 1}$, $y\in\mathbb{Z}_{\geq 0}$ are unknown. Then, there are exactly $L/2$ solutions $(x,y,m)$ if $L$ is even, and (L1)/2(L-1)/2 solutions if $L$ is odd.
@@
By Proposition II.11, m(n)=4(n1)m(n)=4(n-1). Define also
@@
For the second part, by Proposition II.11, we know that the number of powers of simple closed curves of general type with word length exactly $L$ is M(L)=4(L1)M(L)=4(L-1). Hence,
@@ III.1  Characterization
Let $\omega=x_{1}\cdots x_{n}$, where $x_{i}\in\{a,b,a^{-1},b^{-1}\}$ be a reduced word. Consider all of its circular shifts $\{\omega_{i}\}_{i=1}^{n}$ where $\omega_{i}=x_{i}x_{i+1}\cdots x_{n}x_{1}x_{2}\cdots x_{i-1}$ and the cyclic lexicographic ordering induced by $a<b<a^{-1}<b^{-1}$: two words $\alpha=x_{1}\cdots x_{n}$ and $\beta=y_{1}\cdots y_{n}$ satisfy $\alpha<\beta$ if either $x_{1}<y_{1}$, or for some $1<t\leq n$, we have $x_{i}=y_{i}$ for $1\leq i\leq t-1$, and $x_{t}<y_{t}$ under the new ordering obtained by cyclically shifting the original one until it starts with $x_{t-1}^{-1}$.
@@
Figure 4. Surgery on a word of the form $ab^{-1}a^{-i}b\omega$ with $i\geq 2$
@@
Hence, we can restrict to the case where the changes of sign come from subwords of the kind $ab^{-1}a^{-1}b$ (up to renaming $a$ for $a^{-1}$ or $b$ for $b^{-1}$). Hence, take a reduced word $ab^{-1}a^{-1}b\omega$ with $\omega$ a subword such that the exponents on $b$ are $\pm 1$. Note that if $\omega$ starts with an $a$, up to homotopy, one finds the first situation in Figure 5, meaning that the surface is divided into two regions with the startpoint of the curve in one of them and the continuation in the other one. Thus, the curve will intersect at some point the arcs given by the commutator $ab^{-1}a^{-1}b$, and so $i(\omega)\leq i(ab^{-1}a^{-1}b\omega)-1$. Therefore, $\omega$ has to represent a simple word starting with $a$. Moreover, $\omega$ has to be of the form $a^{n_{1}}b\cdots a^{n_{k}}b$, as if it was of the form $a^{n_{1}}b^{-1}\cdots a^{n_{k}}b^{-1}$ the curve $ab^{-1}a^{-1}b\omega$ would contain the subword $ba^{n_{1}}b^{-1}a^{n_{2}}$ and the only such a word with self-intersection one is $bab^{-1}a$ as proved above, or if $\omega=ab^{-1}a^{-1}b$, the word would not be primitive.
@@
Note now that the maps $\alpha_{m},\tilde{\alpha}_{m}$ defined in (II.2) fix the set of conjugacy classes $\{[ab^{-1}a^{-1}b],[(ab^{-1}a^{-1}b)^{-1}]\}$. Hence, considering a reduced circular word of the kind $ab^{-1}a^{-1}b\omega$, such that it has a single self-intersection and such that $\omega=a^{n_{1}}b\cdots a^{n_{k}}b$ with $n_{i}\in\{m,m+1\}$ for some $m\neq 0,1$ represents a primitive simple word, we can apply the reduction $\alpha_{m}(ab^{-1}a^{-1}b\omega)=\omega^{\prime}\alpha_{m}(\omega)$ if $|\{i\mid n_{i}=m\}|>|\{i\mid n_{i}=m+1\}|$ (or with $\tilde{\alpha}_{m}$ otherwise), where $\omega^{\prime}$ is a cyclically reduced representative of the conjugacy classes of the commutators of $a,b$. As we are assuming that $\omega$ represents a simple word, applying finitely many times maps of the family $\alpha_{m}$, $\tilde{\alpha}_{m}$, this reduction acts as $A$ on the necklace $n_{1},\ldots,n_{k}]$ defined in Figure 2(c), and it will reduce to a word of the kind $A^{n}(\omega)=a^{n^{\prime}}b$ up to renaming the generators. Therefore, by inversing $A$ with the appropriate $B_{m}$’s, and given Lemma III.3, after renaming properly the generators such that the representative of the commutator is of the form $ab^{-1}a^{-1}b$, we find that there is a unique permutation of every simple word $\omega$ of the form $a^{n_{1}}b\cdots a^{n_{k}}b$ or $a^{-n_{1}}b\cdots a^{-n_{k}}b$ making $ab^{-1}a^{-1}b\omega$ a self-intersection one curve.
Moreover, applying this reduction to Lemma III.3, one sees that, in order to obtain a self-intersection one curve from a simple one of the form $a^{n_{1}}b\cdots a^{n_{k}}b$ with $n_{i}\in\{m,m+1\}$ for some $m\neq 0,-1$, the commutator $ab^{-1}a^{-1}b$, has to be inserted before a block $a^{n_{i}}b$ such that $n_{i}=m+1$, $n_{i-1}=m$ and maximizing as much as possible the beginning of the chain $\{\sum_{j=1}^{s}|n_{i+\epsilon j}|\}_{s\geq 1}$ with $\epsilon=1$ for positive $m$ and 1-1 for negative (e.g. see Figure 6).
@@
  • $a^{n_{1}}b\cdots a^{n_{k}}b$ where the exponent necklace $n_{1},\dots,n_{k}]$ has $2$-variation.
@@
Let us now move to the case where we have a word of the form $a^{n_{1}}b\cdots a^{n_{k}}b$ with all $n_{i}\in\{m,m+1\}$ for some $m\in\mathbb{Z}_{\geq 1}$. We want to prove that it has self-intersection one if and only if the necklace of positive integers $n_{1},\ldots,n_{k}]$ has $2$-variation.
It is enough to see that in this case there is a $1:1$ correspondence between classes of linking pairs in [CL87] algorithm and essential pairs of blocks in the necklace $n_{1},\ldots,n_{k}]$. We will start by constructing a class of linking pairs in the word given an essential pair of blocks in the associated necklace.